OpenStudy (anonymous):

The perimeter of a rectangle is six times the width. If the length were increased by 4 inches and the width by 7 inches, the perimeter would be 160 inches. What is the width?

OpenStudy (anonymous):

(1)p=6w, and (2) 160=2(l+4)+2(w+7)

OpenStudy (anonymous):

6w=138

OpenStudy (anonymous):

p=6w p=2w+2l so 6w=2w+2l 4w=2l l=2w 160=2(l+4)+2(w+7) plug l=2w 160=2(2w+4)+2(w+7) 160=4w+8+2w+14 138=6w w=23

OpenStudy (anonymous):

160=p+8+14 therefore p=138=6w and w=23 easier :)

OpenStudy (anonymous):

Johnny you are the bestest ever!!!